OpenStudy (anonymous):

An equation in the system is y = x + 1. What is the other equation in the system if the solution is only at (0, 1)? Select one of the options below as your answer: A. x − y = 1 B. 2x + 2y = 4 C. x + y = 1 D. x + 2y = 2

OpenStudy (ivettef365):

just insert the points on the equation to see which one is true example points (0,1) a) x - y = 1 0 -1 = 1 -1 = -1 NOT TRUE incorrect keep trying the others.

OpenStudy (ivettef365):

I meant -1 = 1 not true